X+[x+2] + [x+4] = 87
Answer by checkley79(3341)

X+[x+2] + [x+4] = 87
3X+6=87
3X=87-6
3X=81
X=81/3
X=27 ANS.
PROOF:
27+27+2+27+4=87
87=87